OpenStudy (anonymous):

(integrated) (3x+2)dx/(sqrt 1-x^2)

zepdrix (zepdrix):

Hmmm I think you're going to have to split it into 2 fractions. You can apply a nice easy U substitution to the first fraction, and the second one will require a trig sub it appears :O \[\huge \int\limits\limits \frac{3x+2}{\sqrt{1-x^2}}dx\] \[\huge = 3\int\limits \frac{x}{\sqrt{1-x^2}}dx\quad +\quad 2\int\limits \frac{1}{\sqrt{1-x^2}}dx\]

OpenStudy (anonymous):

for 1/(sqrt 1-x^2) dx = sin^(-1)

zepdrix (zepdrix):

k cool :) so looks like you have 2 of those, and then you have 3 of something else. remember how to do the substitution? :D

OpenStudy (anonymous):

yes i,m already remember .. thank for helping. do you have any method to remeber the inverse trigo.???

zepdrix (zepdrix):

No, arctan is the only one i can remember. the other ones i always have to work out by hand.

OpenStudy (anonymous):

how to derived for arctan ???